OpenStudy (anonymous):

m^3/8(m^-1/4+9m) simplify

OpenStudy (anonymous):

\[\frac{m^3}{8\left(\frac{m^{-1}}{4+9m}\right)} = \frac{m^3(4+9m)}{8m^{-1}} = \frac{m^4(4+9m)}{8}\]

OpenStudy (mertsj):

Why did you post this problem again? And again it is not clear what the problem is.